A source of sound is moving towards a stationary observer with one-ninth the speed of sound. Calculate the ratio of apparent to real frequency
1
Expert's answer
2019-05-09T11:23:39-0400

The apparent frequency is


"f'=f \frac{v}{v-v_a}"

The ratio of apparent to real frequency:

"\frac{f'}{f}= \frac{v}{v-v_a}=\frac{1}{1-\frac{v_a}{v}}"

"\frac{f'}{f}= \frac{1}{1-\frac{1}{9}}=\frac{9}{8}"